Nowlin pipe & steel has projected sales of 72,000 pipes this year, an ordering cost of $6 per order, and carrying costs of $2.40 per pipe.

A. What is the economic ordering quantity?

B. How many orders will be placed during the year?

C. What will be the average inventory?

D. What will be the cost of ordering?

E. What will be the carrying cost?
Answer #1

A.

the economic ordering quantity = ((2*sales*ordering costs) / carrying cost)^(1/2) = (((2*72000*6)/2.40))^(1/2) = 600 units

B.

orders will be placed during the year = 72000/600 = 120

C.

the average inventory = 600/2 = 300

D.

cost of ordering = 6*120 = 720

E.

carrying cost = 2.40*300= 720